OpenStudy (firejay5):

Simplify. Show work and explain how you did it. 3. square root of 3 + sq root of 72 - sq root of 128 + sq root of 108

OpenStudy (firejay5):

\[\sqrt{3} + \sqrt{72} - \sqrt{128} + \sqrt{108}\]

OpenStudy (firejay5):

@jiji501 Are you trying to figure it out???

OpenStudy (anonymous):

umm, wat are the common factors of 72 128 and 108?

OpenStudy (firejay5):

72: 1,2,3,4,6,8,9,12,18,24,36,72 128: 1, 2, 4, 8, 16, 32, 64 and 128 108: 1, 2, 3, 4, 6, 9, 12, 18, 27, 36, 54, 108

OpenStudy (campbell_st):

now identify the square numbers that are factors... e.g. \[\sqrt{72} = \sqrt{36 \times 2} = \sqrt{36} \times \sqrt{2} = 6\sqrt{2}\] use the same process for the other numbers...

OpenStudy (firejay5):

pick the biggest square root

OpenStudy (firejay5):

sq root of 3 is the same correct

OpenStudy (campbell_st):

thats correct... and you should have 2 square rt 3 terms and 2 sq rt 2 terms... just add or subtract then... just the same as like terms in algebra.

OpenStudy (firejay5):

what do I do once I get done with all that

OpenStudy (campbell_st):

thats it,,, what answer did you get..?

OpenStudy (firejay5):

does the square root of 3 have a 1 in front of it

OpenStudy (campbell_st):

nope...its like x its really 1x but we just write x

OpenStudy (firejay5):

but could it

OpenStudy (campbell_st):

it could...

OpenStudy (firejay5):

is the answer 7 square root of 3 - 2 square root of 2

OpenStudy (campbell_st):

thats it... well done...

OpenStudy (firejay5):

I meant to type it in